brettandadrionna:

A business traveler rents a car for $75 a day, and $0.50 for each mile. If the trip will last 3 days and the traveler can spend no more than $400, how many miles can be driven? Explain your reasoning and express the formula as an inequality.

Convert:

75*3= $225 400-225= $175 0.50 *2 = $1 175*2= 350 miles